OpenStudy (anonymous):

I don't even know where to begin when solving this Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold 4.Graph the function. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper

OpenStudy (anonymous):

5.Suppose Sal's total profit on lunch specials for the next month is $1,593. The profit amounts are the same: $2 for each sandwich and $3 for each wrap. In a paragraph of at least three complete sentences, explain how the graphs of the functions for the two months are similar and how they are different

OpenStudy (anonymous):

Anyone???

OpenStudy (baru):

do you recognize this equation: y=mx+c

OpenStudy (zpupster):

baru is giving the hint change the equation from a standard line equation Ax + By =C or in this case 2x + 3y = 1,470 change to slope intercept form y=mx+b to do so divide the equation by 3 and subtract x term from both sides what did you get??